# ChartQA-X: Generating Explanations for Visual Chart Reasoning

### v2

@@ Abstract
The ability to explain complex information from chart images is vital for effective data-driven decision-making. In this work, we address the challenge of generating detailed explanations alongside answering questions about charts. We present ChartQA-X, a comprehensive dataset comprising 30,299 chart samples across four chart types, each paired with contextually relevant questions, answers, and explanations. Explanations are generated and selected based on metrics such as faithfulness, informativeness, coherence, and perplexity. Our human evaluation with 245 participants shows that model-generated explanations in ChartQA-X surpass human-written explanations in accuracy and logic and are comparable in terms of clarity and overall quality. Moreover, models fine-tuned on ChartQA-X show substantial improvements across various metrics, including absolute gains of up to 24.57 points in explanation quality, 18.96 percentage points in question-answering accuracy, and 14.75 percentage points on unseen benchmarks for the same task. By integrating explanatory narratives with answers, our approach enables agents to communicate complex visual information more effectively, improving comprehension and fostering greater trust in the generated responses.
@@ 1 Introduction
Vision language models (VLMs) have made significant progress in addressing complex reasoning tasks [12, 44, 23, 40, 9, 7]. As their capabilities continue to grow, ensuring alignment with human values becomes increasingly critical. However, these models are often viewed as black boxes, producing answers without revealing the reasoning or evidence behind them. Providing explanations helps make their decision-making processes more transparent and interpretable to users. While recent work has highlighted the value of explanations in vision-language reasoning tasks involving general images [38, 45, 31, 28], little research has explored explanation generation for data visualizations. Chart question answering (ChartQA) is an emerging area focused on developing models that can interpret chart data and answer user queries [26, 17, 18, 27]. However, existing ChartQA models produce answers without the explanatory context needed to support user understanding and build trust. This limitation reduces their practical utility, as users often seek not only accurate answers but also clear and interpretable reasoning behind them.
@@
Motivated by this goal, we introduce ChartQA-X, a comprehensive dataset comprising 30,299 chart samples (28,299 train and 2,000 test), each paired with contextually relevant questions, answers, and explanations. Explanations are selected from the outputs of six state-of-the-art vision-language models. ChartQA-X includes a diverse range of chart types, including (1) horizontal bar charts, (2) vertical bar charts, (3) line charts, and (4) pie charts, all derived from the original ChartQA dataset [26]. We assess ChartQA-X through an online human-subject study with 245 participants who rate ChartQA-X explanations as comparable to or better than human-written ones in terms of accuracy, clarity, logic, and overall quality. To establish performance baselines, we fine-tune three state-of-the-art vision-language models, i.e., LLaVA 1.6 [23], Qwen2-VL [40], and InternVL-2.5 [6], on the ChartQA-X dataset and evaluate them in terms of the quality of generating explanations and accuracy of question answering. We further assess their question-answering capabilities on unseen benchmark chart datasets, including DVQA [17], PlotQA [27], and FigureQA [18]. In summary, our contributions are as follows:
  • We introduce ChartQA-X, the largest dataset containing 30,299 chart samples with detailed explanations spanning diverse chart types.
@@ 3.1 Input Data
ChartQA-X is built upon the ChartQA dataset [26], which consists of 18,317 chart images and 30,299 question–answer pairs. It spans four chart types and features a diverse range of descriptive and reasoning questions (Figure 3) with varying levels of complexity. Reasoning questions require calculation using chart data, while descriptive questions involve directly reading values from the chart. Each data point in the ChartQA-X dataset consists of four components: (1) Chart Image ($I$): A visual representation of the data, which may take the form of horizontal bar charts, vertical bar charts, line charts, or pie charts. The image provides visual context for the model, enabling it to interpret the data and incorporate visual attributes, such as color and shape, into its explanations. (2) Data Table ($D$): A structured, tabular representation of the chart’s underlying data, offering precise numerical values and categorical labels. This component complements the visual chart by providing the chart data in a machine-readable format. (3) Question ($Q$): A natural language query about the information presented in the chart. Questions span a variety of types, including trend analysis, comparisons, arithmetic reasoning, and specific data retrieval, serving as the primary prompt for the model. (4) Answer ($A$): The correct response to the question, derived from the chart and its associated table. This is the expected output the model must produce, and the explanation is evaluated based on how well it supports this answer.
@@ 3.2 Explanation Generation
We generate explanations for each question-answer pair using six different VLMs: LLaVA 1.6 [23], Phi-3 [1], CogVLM [41], Deepseek-VL [24], Qwen2-VL [40], and GPT-4o [15]. Each model is prompted to generate explanations in a chain-of-thought reasoning format [42], which breaks down the reasoning process into a sequence of logical steps, making the model’s thought process transparent. This involves analyzing the chart to identify key data points, trends, and relationships, aligning the relevant information with the question, and constructing a coherent narrative that links the data to the final answer. To guide the models in generating such explanations, we design the following prompt:
@@
Generated explanations are limited to four sentences to maintain clarity and brevity.
@@ 3.3 Evaluating and Selecting Explanations
To assess the quality and effectiveness of the explanations, we use the ROSCOE evaluation suite [11], which is specifically designed to score step-by-step reasoning. ROSCOE offers a comprehensive framework that evaluates reasoning quality across nine critical dimensions. In our setup, we define the source context as $s=\{s_{1},...,s_{T}\}$, a sequence of $T$ sentences consisting of the instruction followed by the question. The hypothesis is denoted as $h=\{h_{1},...,h_{N}\}$, a sequence of $N$ reasoning steps, including the final explanation. Table 1 provides the equations for each of the nine ROSCOE metrics.
1. Faithfulness Step (FS) evaluates whether the model misinterprets the problem statement or if the reasoning chain is vague, irrelevant, or misuses information. It measures the alignment ($r{-align}$) between each step of the hypothesis ($h_{i}$) and source context ($s$) and is computed as the mean alignment score across all steps.
@@
Each explanation generated by the six VLMs is evaluated using the nine ROSCOE metrics. Since different models perform best on different subsets of the dataset, we select, for each sample, the explanation with the highest overall score across all metrics. To ensure fair comparison, all metric scores are normalized to a 01010–10 – 1 scale, with 1 indicating the best performance and 0 indicating the worst. For perplexity-based metrics (PS and PC), where lower is better, scores are inverted by subtracting them from 1. We then apply weights between 0 and 1 to each metric to compute an aggregate score. To prioritize alignment between the explanation, question, and answer, we assign higher weights to faithfulness and informativeness metrics and lower weights to others. Based on pilot experiments across various chart types, we set the weights as follows: 0.2 for the faithfulness step, 0.15 each for the faithfulness token and the informativeness chain, 0.1 for self-consistency and source-consistency, and 0.05 each for the perplexity step, perplexity chain, and grammar step. This weighting scheme emphasizes explanation quality and relevance while also considering coherence, fluency, and grammatical accuracy.
[…]
For each sample, we select the explanation with the highest Aggregate Score (AS) among the six models, resulting in a high-quality collection of explanations. Table 2 presents the average ROSCOE scores across all ChartQA samples for each of the six VLMs.
@@
Table 3: Question answering accuracy (%) on different chart and question types (test set only). Best scores are in bold, and second-best scores are underlined.
@@
Table 4: Results of evaluating VLMs on n-gram metrics on the Chart-QA test set. Best scores are in bold, and second-best scores are underlined.

### v4

@@ Abstract
The ability to explain complex information from chart images is vital for effective data-driven decision-making. In this work, we address the challenge of generating detailed explanations alongside answering questions about charts. We present ChartQA-X, a comprehensive dataset comprising 30,799 chart samples across four chart types, each paired with contextually relevant questions, answers, and explanations. Explanations are generated and selected based on metrics such as faithfulness, informativeness, coherence, and perplexity. Our human evaluation with 245 participants shows that model-generated explanations in ChartQA-X surpass human-written explanations in accuracy and logic and are comparable in terms of clarity and overall quality. Moreover, models fine-tuned on ChartQA-X show substantial improvements across various metrics, including absolute gains of up to 24.57 points in explanation quality, 18.96 percentage points in question-answering accuracy, and 14.75 percentage points on unseen benchmarks for the same task. By integrating explanatory narratives with answers, our approach enables agents to convey complex visual information more effectively, improving comprehension and greater trust in the generated responses.
@@ 1 Introduction
Vision language models (VLMs) have made significant progress in addressing complex reasoning tasks [13, 47, 26, 43, 9, 7]. As their capabilities continue to grow, ensuring alignment with human values becomes increasingly critical. However, these models are often viewed as black boxes, producing answers without revealing the reasoning or evidence behind them. Providing explanations helps make their decision-making processes more transparent and interpretable to users. Prior work on visual reasoning for images [11, 49] emphasizes the value of generating answers with explicit rationales to support model predictions. Others [50] leverage large language models for commonsense reasoning, demonstrating improved performance in visual question answering. While recent work has highlighted the value of explanations in vision-language reasoning tasks involving general images [41, 48, 34, 31], little research has explored explanation generation for data visualizations. Chart question answering (ChartQA) is an emerging area focused on developing models that can interpret chart data and answer user queries [29, 18, 19, 30]. However, existing ChartQA models produce answers without the explanatory context needed to support user understanding and build trust. This limitation reduces their practical utility, as users often seek not only accurate answers but also clear and interpretable reasoning behind them.
@@
Motivated by this goal, we introduce ChartQA-X, a comprehensive dataset comprising 30,799 chart samples (28,299 train and 2,500 test), each paired with contextually relevant questions, answers, and explanations. Explanations are selected from the outputs of six state-of-the-art vision-language models. Figure 1 illustrates the ChartQA-X task setup, demonstrating how the dataset enables a model to generate both the answer and an explanation for a given chart-based question. While recent VLMs can generate explanations, their quality varies widely across questions and chart types. Without additional supervision, outputs are often verbose, ungrounded, or inconsistent. To ensure high-quality explanations in ChartQA-X, we simplify the task during dataset creation by supplying full context (e.g., answer, data table) and generating diverse candidate explanations with multiple VLMs. These are then filtered using automatic reasoning metrics to retain only high-quality explanations. ChartQA-X includes a diverse range of chart types, including (1) horizontal bar charts, (2) vertical bar charts, (3) line charts, and (4) pie charts, all derived from the original ChartQA dataset [29]. We assess ChartQA-X through an online human-subject study with 245 participants who rate ChartQA-X explanations as comparable to or better than human-written ones in terms of accuracy, clarity, logic, and overall quality.
By releasing a validated dataset, ChartQA-X enables benchmarking and facilitates progress in VLM explanation generation for chart questions. To establish performance baselines, we fine-tune three state-of-the-art vision-language models, i.e., LLaVA 1.6 [26], Qwen2-VL [43], and InternVL-2.5 [6], on the ChartQA-X dataset and evaluate them in terms of the quality of generating explanations and accuracy of question answering. We further assess their question-answering capabilities on unseen benchmark chart datasets, including DVQA [18], PlotQA [30], and FigureQA [19]. In summary, our contributions are as follows:
  • We introduce ChartQA-X, the largest dataset containing 30,799 chart samples with detailed explanations spanning diverse chart types.
@@ 3.1 Input Data
ChartQA-X is built upon the ChartQA dataset [29], which consists of 18,317 chart images and 30,799 question–answer pairs. It spans four chart types and features a diverse range of descriptive and reasoning questions (Figure 3) with varying levels of complexity. Reasoning questions require calculation using chart data, while descriptive questions involve directly reading values from the chart. Each data point in the ChartQA-X dataset consists of four components: (1) Chart Image ($I$): A visual representation of the data, which may take the form of horizontal bar charts, vertical bar charts, line charts, or pie charts. The image provides visual context for the model, enabling it to interpret the data and incorporate visual attributes, such as color and shape, into its explanations. (2) Data Table ($D$): A structured, tabular representation of the chart’s underlying data, offering precise numerical values and categorical labels. This component complements the visual chart by providing the chart data in a machine-readable format. (3) Question ($Q$): A natural language query about the information presented in the chart. Questions span a variety of types, including trend analysis, comparisons, arithmetic reasoning, and specific data retrieval, serving as the primary prompt for the model. (4) Answer ($A$): The correct response to the question, derived from the chart and its associated table. This is the expected output the model must produce, and the explanation is evaluated based on how well it supports this answer.
@@ 3.2 Explanation Generation
We generate explanations for each question-answer pair using six different VLMs: LLaVA 1.6 [26], Phi-3 [1], CogVLM [44], Deepseek-VL [27], Qwen2-VL [43], and GPT-4o [16]. Each model is prompted to generate explanations in a chain-of-thought reasoning format [45], which breaks down the reasoning process into a sequence of logical steps, making the model’s thought process transparent. This involves analyzing the chart to identify key data points, trends, and relationships, aligning the relevant information with the question, and constructing a coherent narrative that links the data to the final answer. Our final prompt is based on iterative experimentation with several prompt designs to elicit high-quality explanations from VLMs. Our early prompt versions include a fixed set of sub-questions to guide step-wise reasoning, such as “What values are shown?” or “What comparisons are relevant?”. While structured, this approach lacks flexibility and performs poorly across diverse question types. We next attempt an open-ended prompt that instructs the model to generate its own sub-questions before answering them. This produces longer outputs, but often includes irrelevant or redundant steps and conversational phrasing. To improve clarity and relevance, our final prompt is designed following HCI literature to emphasize concise reasoning with visual references like human explanations [21], limit responses to four sentences, and avoid conversational language. Specifically, we design the following prompt:
@@
We note that explanation generation during dataset construction is not equivalent to explanation generation at inference time. During this phase, we provide models with full context (the chart, question, answer, and data table), which significantly lowers task complexity. Furthermore, to reduce model-specific bias, we sample explanations from six different VLMs. As a result, the final dataset is composed of explanations not tied to a single model’s reasoning style. This strategy mitigates circularity and enables models to learn from superior or complementary reasoning patterns present across models.
@@ 3.3 Evaluating and Selecting Explanations
To assess the quality and effectiveness of the explanations, we use the ROSCOE evaluation suite [12], which is specifically designed to score step-by-step reasoning. ROSCOE offers a comprehensive framework that evaluates reasoning quality across nine critical dimensions. In our setup, we define the source context as $s=\{s_{1},...,s_{T}\}$, a sequence of $T$ sentences consisting of the instruction followed by the question. The hypothesis is denoted as $h=\{h_{1},...,h_{N}\}$, a sequence of $N$ reasoning steps, including the final explanation. Reasoning Alignment (ralignr{-align}) measures the degree to which each step in a multi-step explanation aligns with the source context (i.e., question). Alignment for each step is computed as:
$r\text{-align}(h_{i}\to s)=\frac{[1+\max_{j=1}^{T}(\cos(h_{i},s_{j})]}{2}$ (1)
yielding a vector $\in[0,1]^{N}$. This reflects the semantic closeness of each reasoning step to the source. Contradiction Probability ($p_{\text{contradicts}}$) is estimated using a Natural Language Inference (NLI) model, in this case a fine-tuned DeBERTa-v3 [23], that classifies explanation-source pairs into entailment, neutral, or contradiction. The $p_{\text{contradicts}}$ score is the model’s confidence that the explanation contradicts the source, serving as a proxy for logical inconsistency. English grammatical acceptability $p_{gram}$ is scored by a classifier model [22]. Table 1 provides the equations for each of the nine ROSCOE metrics.
1. Faithfulness Step (FS) evaluates whether the model misinterprets the problem statement or if the reasoning chain is vague, irrelevant, or misuses information. It measures the alignment (ralignr{-align}) between each step of the hypothesis ($h_{i}$) and source context ($s$) and is computed as the mean alignment score across all steps.
@@
Each explanation generated by the six VLMs is evaluated using the nine ROSCOE metrics. Since different models perform best on different subsets of the dataset, we select, for each sample, the explanation with the highest overall score across all metrics. To ensure fair comparison, all metric scores are normalized to a 0-11 scale, with 1 indicating the best performance and 0 indicating the worst. For perplexity-based metrics (PS and PC), where lower is better, scores are inverted by subtracting them from 1. We then apply weights between 0 and 1 to each metric to compute an aggregate score. To prioritize alignment between the explanation, question, and answer, we assign higher weights to faithfulness and informativeness metrics and lower weights to others. We set the weights as follows: 0.2 for the faithfulness step, 0.15 each for the faithfulness token and the informativeness chain, 0.1 for self-consistency and source-consistency, and 0.05 each for the perplexity step, perplexity chain, and grammar step. This weighting scheme emphasizes explanation quality and relevance while also considering coherence, fluency, and grammatical accuracy. To evaluate the stability of this approach, we conduct a sensitivity analysis using a variety of alternative weight configurations, including uniform, random, and metric-focused schemes. The results, presented in Appendix A.3, show that while weighting affects selection, our choices align well with other meaningful configurations.
We compute an Aggregate Score (AS) for each explanation as a weighted sum of all normalized metric scores. For each sample, we select the explanation with the highest Aggregate Score (AS) among the six models, resulting in a high-quality collection of explanations. Table 2 presents the average ROSCOE scores across all ChartQA samples for each of the six VLMs.
@@
Table 3: Question answering accuracy (%), without data table in the input, calculated for different chart and question types (test set only). Best scores are in bold, and second-best scores are underlined.
Models Chart Types Question Types Overall Horizontal Bar Vertical Bar Pie Line Descriptive Reasoning State-of-the-Art Models GPT-4o [16] 76.45 76.43 84.88 70.07 75.74 76.31 76.64 Deepseek-VL [27] 59.07 62.26 56.98 52.37 68.20 48.87 57.96 CogVLM [44] 60.74 61.74 65.12 50.62 63.68 55.40 59.55 Phi-3 [1] 74.65 82.96 74.42 72.32 84.47 70.56 76.56 Models Fine-tuned on ChartQA-X LLaVA-1.6 [26] 48.39 55.74 53.49 42.14 57.40 43.73 50.15 + ChartQA-X 60.36 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}+11.97}$ 62.70 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}+6.96}$ 63.95 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}+10.46}$ 52.87 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}+10.73}$ 67.83 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}+10.43}$ 51.83 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}+8.10}$ 59.92 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}+9.77}$ Qwen2-VL [43] 72.90 77.61 72.49 65.83 82.02 62.94 72.30 + ChartQA-X 74.90 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}+2.0}$ 80.95 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}+3.34}$ 77.91 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}+5.42}$ 67.83 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}+2.0}$ 85.72 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}+3.70}$ 66.20 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}+3.26}$ 75.59 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}+3.29}$ InternVL-2.5 [7] 75.42 80.35 76.16 71.57 85.65 67.07 76.04 + ChartQA-X 79.28 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}+3.86}$ 83.74 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}+3.39}$ 83.72 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}+7.56}$ 72.82 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}+1.25}$ 87.35 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}+1.70}$ 72.65 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}+5.58}$ 79.93 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}+3.89}$
Table 4: Question answering accuracy (%), with data table included in the input, calculated for different chart and question types (test set only). Best scores are in bold, and second-best scores are underlined.
Models FS FT IS IC SRC SFC PS PC GS AS State-of-the-Art Models Phi-3 [1] 0.865 0.939 0.873 0.925 0.678 0.792 0.995 0.947 0.946 0.875 CogVLM [44] 0.873 0.938 0.887 0.919 0.471 0.714 0.997 0.997 0.937 0.851 Deepseek-VL [27] 0.871 0.939 0.884 0.921 0.311 0.722 0.996 0.928 0.928 0.832 GPT-4o [16] 0.882 0.941 0.892 0.916 0.803 0.850 0.972 0.972 0.937 0.899 Models Fine-tuned on ChartQA-X InternVL-2.5 [7] 0.862 0.939 0.868 0.918 0.715 0.722 0.995 0.959 0.951 0.870     + ChartQA-X 0.894 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}+0.032}$ 0.947 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}+0.008}$ 0.877 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}+0.009}$ 0.922 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}+0.004}$ 0.797 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}+0.082}$ 0.853 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}+0.131}$ 0.994 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}-0.001}$ 0.963 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}+0.004}$ 0.958 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}+0.007}$ 0.901 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}+0.031}$ LLaVA 1.6 [26] 0.884 0.945 0.877 0.901 0.515 0.858 0.989 0.991 0.941 0.869     + ChartQA-X 0.928 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}+0.044}$ 0.949 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}+0.004}$ 0.894 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}+0.017}$ 0.918 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}+0.017}$ 0.609 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}+0.094}$ 0.920 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}+0.062}$ 0.997 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}+0.008}$ 0.996 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}+0.005}$ 0.980 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}+0.039}$ 0.901 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}+0.032}$ Qwen2-VL [43] 0.868 0.941 0.868 0.918 0.674 0.871 0.983 0.995 0.964 0.884     + ChartQA-X 0.907 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}+0.039}$ 0.953 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}+0.012}$ 0.884 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}+0.016}$ 0.963 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}+0.045}$ 0.902 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}+0.228}$ 0.918 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}+0.047}$ 0.997 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}+0.014}$ 0.997 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}+0.002}$ 0.970 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}+0.006}$ 0.932 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}+0.048}$
Table 5: ROSCOE scores on the ChartQA-X test set without data table in the input. Best scores are in bold, and second-best scores are underlined. FS: Faithfulness Step, FT: Faithfulness Token, IS: Informativeness Step, IC: Informativeness Chain, SRC: Source-Consistency, SFC: Self-Consistency, PS: Perplexity Step, PC: Perplexity Chain, GS: Grammar Step, and AS: Aggregate Score.
@@
Table 6: Evaluation results on n-gram metrics for the Chart-QA test set. Best scores are in bold, and second-best scores are underlined.